A 110 kg hockey player moving at 4.5 m/s runs into a 90 kg hockey player that is moving at 2.5 m/s. After the collision, the 110 kg player is moving at 3.25 m/s, what is the new speed of the 90 kg hockey player? 


1
Expert's answer
2021-03-29T18:54:23-0400

Let's apply the law of conservation of momentum:


"m_1v_{1i}+m_2v_{2i}=m_1v_{1f}+m_2v_{2f},""v_{2f}=\dfrac{m_1v_{1i}+m_2v_{2i}-m_1v_{1f}}{m_2},""v_{2f}=\dfrac{110\ kg\cdot4.5\ \dfrac{m}{s}+90\ kg\cdot2.5\ \dfrac{m}{s}-110\ kg\cdot3.25\ \dfrac{m}{s}}{90\ kg}=4.03\ \dfrac{m}{s}."